OpenStudy (anonymous):

which equation represents the line that passes through the point (1,5) and has a slope of -2?

OpenStudy (anonymous):

The coordinate point is (x,y). A linear equation is y=mx+b where m=slope. You would plug in the values for x and y in the formula. \[y=mx+b\]\[5=-2(1)+b\]Then solve for b. After that, you should have the equation that you're looking for.

OpenStudy (anonymous):

1. y=-2x+7 2. y=-2x+11 3. y= 2x-9 4. y= 2x+3 Which one is it?

OpenStudy (anonymous):

After you solve for b, you should get b=7. You plug that into y=mx+b. So it will be y=-2x+7

OpenStudy (anonymous):

thanks! its correct!

OpenStudy (anonymous):

:D